### v3

# Hypercube minor-universality

Abstract.
A graph $G$ is $m$-minor-universal if every graph with at most $m$ edges (and no isolated vertices) is a minor of $G$. We prove that the $d$-dimensional hypercube, $Q_{d}$, is $\Omega\left(\frac{2^{d}}{d}\right)$-minor-universal, and that there exists an absolute constant $C>0$ such that $Q_{d}$ is not $\frac{C2^{d}}{\sqrt{d}}$-minor-universal. Similar results are obtained in a more generalized setting, where we bound the size of minors in a product of finite connected graphs. A key component of our proof is the following claim regarding the decomposition of a permutation of a box into simpler, one-dimensional permutations: Let $n_{1},\dots,n_{d}$ be positive integers, and define $X:=[n_{1}]\times\dots\times[n_{d}]$. We prove that every permutation $\sigma:X\to X$ can be expressed as $\sigma=\sigma_{1}\circ\dots\circ\sigma_{2d-1}$, where each $\sigma_{i}$ is a one-dimensional permutation, meaning it fixes all coordinates except possibly one. We discuss future directions and pose open problems.
1. Introduction
By contraction of an edge $uv$ in a graph $G$, we mean identification of $u$ and $v$, i.e., replacement of $u$ and $v$ by a new vertex $w$ adjacent to all of the neighbors of $u$ and $v$. A graph $H$ is a minor of a graph $G$ if by repeatedly deleting vertices and edges from $G$, as well as contracting edges, one can obtain a graph isomorphic to $H$. Observe that the minor relation is transitive, that is, if $F\leq H$ and $H\leq G$ then $F\leq G$.
Definition 1.1.
A graph $G$ is $\bm{m}$-minor-universal if $H\leq G$ for every graph $H$ with at most $m$ edges and with no isolated vertices. The minor-universality of $G$, $m(G)$, is the maximal $m\in\mathbb{N}\cup\{0\}$ such that $G$ is $m$-minor-universal.
We study the existence and non-existence of minors using combinatorial embeddings. Determining which graphs can be embedded in graph $G$ via combinatorial embeddings is nearly equivalent to identifying the minors of $G$. Section 2 formally defines these embeddings and clarifies their precise relationship to minors.
Combinatorial embeddings are a combinatorial adaptation of thick embeddings. The study of thick embeddings traces back to the 1960s to the seminal work of Kolmogorov and Barzdin [KB93]. As part of their efforts to model physical networks, such as neural networks in the brain or logical networks in computers, they asked how tightly can such networks be realized in 3 dimensional space. They showed that any graph with $m$ edges (and no isolated vertices) can be embedded in a Euclidean ball of volume $O(m^{3/2})$ using a 1-thick embedding. Moreover, to prove this is the best possible they essentially gave the first construction of expander graphs. Even though they worked with the continuous notion of thick embeddings, one can transfer their result to the discrete settings, and get that the 3-dimensional ball of radius $cm$ in $\mathbb{Z}^{3}$ is $m^{2}$-minor-universal for large enough $c>0$.
Since the seminal work of Kolmogorov and Barzdin, thick embeddings were further studied in [GG12], [BH21], [Ka24] in the context of thick embeddings of graphs and higher dimensional simplicial complexes into various spaces, not necessarily euclidean.
We highlight that from a combinatorial perspective, embedding minors into graphs, and in particular establishing which graphs are $m$-minor-universal was studied in computer science and combinatorics. In [KR96] Kleinberg and Rubinfeld  proved that an $\alpha$-expander with $n$ vertices and maximum degree $\Delta$ is $\Omega\left(\frac{n}{(\log n)^{\kappa}}\right)$-minor-universal for some $\kappa=\kappa(\alpha,\Delta)>1$. Krivelevich and Nenadov [Kr19] showed that if $G$ is an $\alpha$-expander, then $m(G)=\Omega\left(\frac{n}{\log n}\right)$, and if $G$ also has logarithmic girth, then $m(G)=\Theta\left(\frac{n}{\log n}\right)$.
There is also a significant body of work on clique minors in general graphs and in random graphs. This line of work is motivated by Hadwiger’s conjecture, e.g., [BCE80, FKO08]. Clique minors in graphs without sparse cuts and in expander graphs have been studied in [KN21] and [KS06], respectively. Results in [EKK23] discuss clique minors in the hypercube itself, and [CS07] examine clique minors in product graphs.
The $d$-dimensional hypercube graph $Q_{d}$ is the graph with $d$-bits binary strings (the set $\{0,1\}^{d}$) as vertices and two binary strings are neighbors if they are of hamming distance $1$. Our main theorem is:
Theorem A (hypercube minor-universality).
$m(Q_{d})=\Omega\left(\frac{2^{d}}{d}\right)$ and $m(Q_{d})=O\left(\frac{2^{d}}{\sqrt{d}}\right)$.
In other words, there are absolute constants $C,C^{\prime}>0$ such that every graph with $\leq C\frac{2^{d}}{d}$ edges (and no isolated vertices) is a minor of $Q_{d}$, while there exists a graph with $\leq C^{\prime}\frac{2^{d}}{\sqrt{d}}$ edges (and no isolated vertices) that is not a minor of $Q_{d}$.
We highlight that this improves on the result of Kleinberg and Rubinfeld [KR96] which guarantees a lower bound $\Omega\left(\frac{2^{d}}{d^{\kappa}}\right)$ for some $\kappa>1$. We also remark that the result of Krivelevich and Nenadov [Kr19] gives a lower bound of $m(Q_{d})=\Omega\left(\frac{2^{d}}{d^{2}}\right)$, hence our lowerbound is a $d$ factor sharper. This follows from the fact that $Q_{d}$ is a $\Theta(1/\sqrt{d})$-expander and the constant in their lower bound depends quadratically on the expansion $\alpha$.
Question 1.
The upper and lower bounds in theorem A do not match; there is a factor of $\sqrt{d}$ between them. Is either of these bounds tight, and if so, which one?
Question 2.
Given $m\in\mathbb{N}$, what are the graphs with minimal number of edges that are $m$-minor-universal?
Question 3.
What is the minor-universality of these two variants of the hypercube: the cube-connected cycles [^1] https://en.wikipedia.org/wiki/Cube-connected_cycles. and the randomly twisted hypercube [BDGZ24]? Additionally, what is the minor-university of the Cayley graph of $S_{n}$ with respect to all transpositions?
We tried to slightly generalize the hypercube theorem and found an extension: The cube $Q_{d}$ can be expressed as a product of graphs, $Q_{d}=Q_{1}\Box\ldots\Box Q_{1}$. [^2] This is the Cartesian product (also known as the box product), which is properly defined in Section 3. This naturally leads to the question of whether a similar result holds for a product of arbitrary sequence of graphs. As demonstrated by our next theorem, such a generalization is indeed achievable.
Theorem B (product graph minor-universality).
Let $k,n\in\mathbb{N}$, and let $G_{1},\dots,G_{n}$ be a sequence of graphs, each one with $\leq k$ vertices.
  1. (1)
    If the graphs $G_{i}$ are connected, then for any pair of connected graphs $H_{4}$ (with 4 vertices) and $H_{k}$ (with $k$ vertices), the graph
    $H_{4}\Box C_{6n-2}\Box H_{k}\Box G_{1}\Box\cdots\Box G_{n}$
    is $\frac{1}{16}|V(G_{1})|\cdots|V(G_{n})|$-minor-universal.
  2. (2)
    There exists a constant $C=C(k)$ such that the graph $G_{1}\Box\cdots\Box G_{n}$ is not $\frac{Ck^{n}}{\sqrt{n}}$-minor-universal.
This theorem generalizes theorem A. The first part extends the lower bound, the result that $Q_{d}$ is $\Omega\left(\frac{2^{d}}{d}\right)$-minor-universal. Although this connection may not be immediately obvious, and the inclusion of the unusual graphs $H_{4},H_{k}$ and the cycle $C_{6n-2}$ might seem somewhat unexpected, all of this is explained in detail in Section 3. The second part of the theorem generalizes the upper bound, which is more straightforward to recognize.
Note that if we treat $n$ as a constant (denoted $d-1$) and $k$ as a variable (denoted $n$), we observe that for $H_{n}=G_{1}=\cdots=G_{d-1}:=P_{n}$ (the path with $n$ vertices), the theorem implies that $F\Box B_{n}$ is $\Omega(n^{d-1})$-minor-universal where $F$ is a fixed finite graph, and $B_{n}$ represents the ball of radius $n$ in $\mathbb{Z}^{d}$. This result is similar to the result stating that for every bounded-degree finite graph $\Gamma$, there exists a $1$-thick embedding into $\mathbb{R}^{d}$ (for $d\geq 3$) with volume $O(|V\Gamma|^{\frac{d}{d-1}})$ [KB93], [Gu16], [BH21]. [^3] One might wonder if [BH21, Theorem 1.3] implies theorem A. While [BH21, Theorem 1.3] demonstrates the embedding of graphs into small balls in $\mathbb{R}^{d}$, it does not guarantee embeddings within the hypercube. Moreover, even if [BH21, Theorem 1.3] could be applied to the hypercube, its lower bound on the minor-universality is only $o\left(\frac{2^{d}}{d}\right)$.
This raises the question of whether taking the product of a graph with a fixed finite graph significantly impacts its minor-universality. Consider the following question:
Question 4.
Give sufficient conditions for a sequence of finite graphs $F,G_{1},G_{2},\ldots$, to satisfy $\forall n:m(F\Box G_{n})\leq Cm(G_{n})$ for some $C<\infty$.
For example, this is false when $G_{n}$ is a $n\times n$ planar grid and $F$ is a single edge. In this case, $m(G_{n})$ is bounded, while $m(F\Box G_{n})$ grows unbounded.
We also believe the following question is an attractive generalization of our main result:
Question 5.
Let $G$ be a connected transitive graph of diameter $D$, is $Q_{1}\Box G$ necessarily $\Omega(\frac{|V(G)|}{D})$-minor-universal? Which connected transitive graphs of diameter $D$ are $\Omega(\frac{|V(G)|}{D})$-minor-universal?
A key part of our proof is the following proposition:
Proposition 1.2 (box permutation decomposition).
Let $d\in\mathbb{N}$ and $n_{1},\dots,n_{d}\in\mathbb{N}$. Define $X:=[n_{1}]\times\dots\times[n_{d}]$. Every permutation $\sigma:X\to X$ can be expressed as
$\sigma=\sigma_{1}\circ\dots\circ\sigma_{2d-1},$
where $\sigma_{1},\dots,\sigma_{2d-1}$ are one-dimensional permutations. Moreover, if $n_{1},\dots,n_{d}\geq 2$, there exists a permutation $\sigma:X\to X$ such that any decomposition of $\sigma$ into one-dimensional permutations requires at least $12d-$ 1 permutations.
A permutation $\sigma:X\to X$ is called one-dimensional if it affects only one coordinate. Formally, $\sigma$ is one-dimensional if there exists $i$ such that for every $x\in X$ the points $x$ and $\sigma(x)$ agree on all coordinate except maybe the $i$-th coordinate.
1.1. Notations
  1. $\bullet$
    We define $[n]:=\{1,\dots,n\}$.
  2. $\bullet$
    $\log n:=\log_{2}n$.
  3. $\bullet$
    All the graphs will be finite and simple (without multi-edges and self loops).
  4. $\bullet$
    Matching graph is graph with maximal degree at most 1.
  5. $\bullet$
    If $l$ is some finite sequence then $l[0]$ is the first element, and $l[-1]$ is the last one.
1.2. Paper overview
Section 2 introduces the concept of combinatorial embedding, a notion closely related to minors. Instead of working directly with minors, we use combinatorial embeddings to prove theorems A and B. In Section 3, we establish the first part of theorem B (which implies the lower bound of theorem A), treating proposition 1.2 as a black box. The proof of proposition 1.2 is provided in Section 4. Finally, in Section 5, we complete the proof of theorem A by establishing the upper bound. This section also includes the proof of the second part of theorem B, which is similar but requires more extensive computations.
1.3. Acknowledgments
Thanks to Michael Krivelevich for providing us with relevant references. E.T. appreciates the support of the Adams Fellowship Program of the Israel Academy of Sciences and Humanities.
2. Combinatorial embedding
Definition 2.1 (combinatorial embedding).
Let $G$ and $Y$ be graphs. A combinatorial embedding of $G=(V,E)$ into $Y$ is a function
$f:V\cup E\to\bigcup_{n\in\mathbb{N}}(VY)^{n}$
such that the image of each vertex is a vertex, and the image of an edge $e=uv\in EG$, denoted by $f_{e}$, is a walk with endpoints $fu,fv$. We call this walk a road. Moreover the followings hold:
  1. (1)
    $f|_{V}$ is injective.
  2. (2)
    ”Roads do not intersect” — for any two non-adjacent edges $e,t\in E$, we have $f_{e}\cap f_{t}=\emptyset.$
  3. (3)
    ”Vertices are isolated from roads” — for any edge $e\in E$ and any vertex $v\notin e$, we have $fv\notin f_{e}.$
Combinatorial embeddings and minors are closely related concepts, as shown in the following two lemmas:
Lemma 2.2.
Let $H$ and $G$ be graphs. If $H\leq G$, then there exists a combinatorial embedding of $H$ into $G$.
To prove this, we begin with the following definition: A model of a graph $H$ in a graph $G$ is a function $\mu$ that assigns to each vertex of $H$ a vertex-disjoint connected subgraph of $G$, such that if $uv\in E(H)$, then some edge in $G$ connects a vertex in $\mu(u)$ to a vertex in $\mu(v)$.
It is known that there exists a model of a graph $H$ in a graph $G$ if and only if $H\leq G$ [N17, Lemma 1.2]. Thus, it suffices to prove that if there is a model of a graph $H$ in $G$, then there exists a combinatorial embedding of $H$ into $G$, indeed:
Proof.
Assume that $\mu$ is a model of $H$ in $G$. We aim to construct a combinatorial embedding $f:H\to G$. For each vertex $v\in V(H)$, define $fv$ as some vertex within $\mu(v)$. For each edge $uv\in E(H)$, there exists an edge $e\in E(G)$ that connects a vertex in $\mu(u)$ to a vertex in $\mu(v)$. Since $\mu(u)$ and $\mu(v)$ are connected, there is a walk from $fu$ to $fv$ that remains entirely within $V(\mu(u))\cup V(\mu(v))$. Define $f_{e}$ as this walk. This construction defines a combinatorial embedding, which is straightforward to verify. ∎
We now consider the other direction, namely, whether the existence of a combinatorial embedding of $H$ into $G$ implies that $H\leq G$. Unfortunately, this is not true, as shown in the image.
Figure 1. There is a combinatorial embedding of the triangle into the star graph, but the triangle is not a minor of the star graph.
However, by slightly modifying $H$, we can establish a related result that holds. The issue with converting a combinatorial embedding into a model is that, for two different edges sharing a common vertex, there is no guarantee about how their corresponding roads will behave. They may overlap excessively, as seen in the triangle example, which is problematic. To address this, we introduce the following definition to separate the roads and ensure that all edges are effectively non-adjacent.
Definition 2.3.
For a graph $G=(V,E)$, the simple subdivision of $G$ is the graph $G^{\prime}=(V^{\prime},E^{\prime})$ where
$V^{\prime}:=V\cup\{(v,u)\mid v\sim u\}$
and
$E^{\prime}:=\{\{v,(v,u)\}\mid v\sim u\}\cup\{\{(v,u),(u,v)\}\mid u\sim v\}.$
In simple words, the simple subdivision of a graph is the graph obtained by taking the original graph and adding two new vertices ”inside” each edge. Each such new vertex is related to one of the endpoints of the edge.
We can now state the following lemma:
Lemma 2.4.
Let $H$ and $G$ be graphs. If there is a combinatorial embedding of the simple subdivision of $H$ into $G$, then $H\leq G$.
Proof.
Let $H^{\prime}$ be the simple subdivision of $H$, and let $f:H^{\prime}\to G$ be a combinatorial embedding. To prove $H\leq G$, we will construct a model $\mu$ of $H$ in $G$.
For each vertex $v\in V(H)$, define its star set as
$S_{v}:=\bigcup_{v\sim u}f_{\{v,(v,u)\}}.$
Notice that these star sets are pairwise disjoint since $f$ is a combinatorial embedding.
For the edge $e^{\prime}=\{(u,v),(v,u)\}\in E(H^{\prime})$, consider the road $f_{e^{\prime}}=(x_{1},\ldots,x_{n})$ where each $x_{i}$ is a vertex in $G$, $x_{1}=f((u,v))$, and $x_{n}=f((v,u))$. Observe that the road starts at a point in $S_{u}$ and ends at a point in $S_{v}$. Along the way, it may enter and exit $S_{u}$ and $S_{v}$ arbitrarily. However, we can restrict a path segment $(x_{i},\ldots,x_{j})$ so that $x_{i}\in S_{u}$, $x_{j}\in S_{v}$, and $x_{k}\in(S_{u}\cup S_{v})^{C}$ for any $i<k<j$. Define the sets $R_{u,v}:=\{x_{i}\}$ and $R_{v,u}:=\{x_{i+1},\ldots,x_{j}\}$ and note that
($\ast$) $R_{u,v}\cap S_{v}=\emptyset\quad\text{and}\quad R_{v,u}\cap S_{u}=\emptyset.$
For each vertex $v\in V(H)$, define $\mu(v)$ as the subgraph of $G$ induced by the vertex set
$V_{v}:=S_{v}\cup\bigcup_{w\sim v}R_{v,w}.$
The subgraph $\mu(v)$ is connected because every vertex in it is connected to $fv$. Additionally, for distinct vertices $u,v\in V(H)$, we have $V(\mu(u))\cap V(\mu(v))=V_{u}\cap V_{v}=\emptyset$, as shown by the following:
  • $S_{u}\cap S_{v}=\emptyset$.
  • For $w\sim v$, we have $S_{u}\cap R_{v,w}=\emptyset$:
    • If $w\neq u$, then $R_{v,w}\subseteq f_{\{(v,w),(w,v)\}}$, a road that is disjoint from all roads forming $S_{u}$.
    • If $w=u$, then $S_{u}\cap R_{v,w}=S_{u}\cap R_{v,u}=\emptyset$ by ($\ast$2).
  • For $w\sim u$ and $w^{\prime}\sim v$, $R_{u,w}\cap R_{v,w^{\prime}}=\emptyset$:
    • If $\{u,w\}\neq\{v,w^{\prime}\}$, then $R_{u,w}\subseteq f_{\{(u,w),(w,u)\}}$ and $R_{v,w^{\prime}}\subseteq f_{\{(v,w^{\prime}),(w^{\prime},v)\}}$, which are disjoint roads.
    • If $\{u,w\}=\{v,w^{\prime}\}$, then $w=v$ and $w^{\prime}=u$, so $R_{u,w}\cap R_{v,w^{\prime}}=R_{u,v}\cap R_{v,u}=\emptyset$ by definition.
Thus, $\mu$ is a model of $H$ in $G$. ∎
3. Lower bound: minor-universality
Here we prove the first part of theorem B, which implies the lower bound of theorem A - the assertion that $Q_{d}$ is $\Omega\left(\frac{2^{d}}{d}\right)$-minor-universal. We present here the proof of the generalized theorem, as it not only generalizes the original result but also sheds light on the real essence of the proof, offering a more natural and comprehensive perspective.
Recall that the Cartesian product of two graphs, $G$ and $H$, is the graph $G\Box H$ with vertex set $V(G)\times V(H)$, where vertices $(g,h)$ and $(g^{\prime},h^{\prime})$ are adjacent if either $g=g^{\prime},h\sim h^{\prime}$ or $g\sim g^{\prime},h=h^{\prime}$. Observe that if $H\leq G$, then $H\Box L\leq G\Box L$ for any graph $L$.
For the rest of this section, fix $k,n\in\mathbb{N}$, and let $H_{4},H_{k},G_{1},\ldots,G_{n}$ be a sequence of connected graphs such that $|V(H_{4})|=4$, $|V(H_{k})|=k$, and $|V(G_{i})|\leq k$ for all $i$. Define $V_{i}:=V(G_{i})$, and label the vertices of each of these graphs by $\{0,1,\ldots\}$. We also consider the cycle graph $C_{6n-2}$, and also labeled its vertices by $\{0,1,\ldots\}$.
Theorem 3.1 (lower bound of theorem B).
The graph
$H_{4}\Box C_{6n-2}\Box H_{k}\Box G_{1}\Box\ldots\Box G_{n}$
is $\frac{1}{16}|V_{1}|\cdots|V_{n}|$-minor-universal.
3.1. Why this theorem is a generalization?
Why is this theorem a generalization of the claim that $Q_{d}$ is $\Omega\left(\frac{2^{d}}{d}\right)$-minor-universal? Let $d\in\mathbb{N}$, and set $n:=d-\lfloor\log d\rfloor-8$. Applying the theorem with $k=2$, $H_{4}=Q_{2}$, $H_{k}=Q_{1}$, and $G_{i}=Q_{1}$, we conclude that the graph $Q_{2}\Box C_{6n-2}\Box Q_{n+1}$ is $\Omega(2^{n})$-minor-universal.
Since $C_{6n-2}\leq Q_{\lceil\log(6n-2)\rceil}$, it follows that $Q_{n+\lceil\log(6n-2)\rceil+3}$ is $\Omega(2^{n})$-minor-universal. Compute:
$n+\lceil\log(6n-2)\rceil+3<n+\log n+7<d-\log d-7+\log d+7=d.$
Thus, also $Q_{d}$ is $\Omega(2^{n})$-minor-universal. Additionally, since $2^{n}\geq 2^{-8}\frac{2^{d}}{d}$, we conclude that $Q_{d}$ is $\Omega\left(\frac{2^{d}}{d}\right)$-minor-universal.
3.2. Proof plan
Before presenting the proof, we introduce additional notation to improve readability:
$P_{0}$ $:=G_{1}\Box\cdots\Box G_{n},$
$P_{1}$ $:=H_{k}\Box P_{0},$
$P_{2}$ $:=C_{6n-2}\Box P_{1},$
$P_{3}$ $:=H_{4}\Box P_{2}.$
Our goal is to show that $P_{3}$ has a lot of minors. By Lemma 2.4, it suffices to show that a lot of graphs can be embedded into $P_{3}$. The original incentive was that a lot of graphs can be embedded (up to small errors) inside $P_{0}$ the product of a sequence of connected graphs. The addition of $H_{4}\Box C_{6n-2}\Box H_{k}$ is an artifact that enable our embeddings. This artifact acts as a type of ”coloring” that separates roads of the embedding.
The proof plan proceeds as follows: we start with a graph with a small number of edges that needs to be embedded into $P_{3}$. After reducing this problem to the problem of embedding a graph with a maximum degree $\leq 3$ and essentially same number of edges, we solve the reduced problem using the following three-step solution:
  1. (1)
    Piecewise embedding of matching graph:We introduce a weaker notion of embedding, called a piecewise embedding. To enable the construction of such an embedding, we add $H_{k}$ in front of $P_{0}$. Our goal is to find a piecewise embedding of a matching graph within $P_{1}=H_{k}\Box P_{0}$, where the vertex positions are predetermined. This is achievable—using the box permutation decomposition, we construct such a $(6n-3)$-piecewise embedding.
  2. (2)
    From piecewise embedding to combinatorial embedding: Adding $C_{6n-2}$ in front of $P_{1}$, we convert a $(6n-3)$-piecewise embedding inside $P_{1}$ to a combinatorial embedding inside $P_{2}=C_{6n-2}\Box P_{1}$.
  3. (3)
    From matching graph to graph with maximal degree $\leq 3$: Adding $H_{4}$ in front of $P_{2}$, we use the ability to embed matching graphs with given vertex positions into $P_{2}$ to construct a combinatorial embedding of graph with maximal degree $\leq 3$ into $P_{3}=H_{4}\Box P_{2}$.
Now we present the proof itself:
3.3. Reduction
It’s enough to prove
Theorem 3.2.
Every graph with maximal degree $\leq 3$ and $\leq\frac{1}{4}|V_{1}|\cdots|V_{n}|$ vertices is a minor of $P_{3}$.
Given a graph $\Gamma$ with no isolated vertices and $m\leq\frac{1}{16}|V_{1}|\cdots|V_{n}|$ edges, we construct a new graph $\Gamma^{\prime}$ by replacing each vertex $v$ of $\Gamma$ with a binary tree having $\deg(v)$ leaves and at most $2\deg(v)$ vertices in total. The binary trees are connected such that if $u$ and $v$ are neighbors in $\Gamma$, an edge is added between the leaf corresponding to $u$ in the tree of $v$ and the leaf corresponding to $v$ in the tree of $u$.
Note that $\Gamma$ is a minor of $\Gamma^{\prime}$. Furthermore, $\Gamma^{\prime}$ has at most $\sum_{v\in V(H)}2\deg(v)=4m\leq\frac{1}{4}|V_{1}|\cdots|V_{n}|$ vertices and a maximum degree of at most $3$. By the theorem, this implies $\Gamma^{\prime}\leq P_{3}$, and therefore $\Gamma\leq P_{3}$. In conclusion, $P_{3}$ is $\frac{1}{16}|V_{1}|\cdots|V_{n}|$-minor-universal, as we wanted.
To prove theorem 3.2, we proceed by the planed three-steps solution:
3.4. Step 1: Piecewise embedding of matching graph
Definition 3.3 (piecewise embedding).
For graphs $\Gamma$ and $Y$ and $N\in\mathbb{N}$, an $\bm{N}$-piecewise embedding from $\Gamma$ to $Y$ is a pair
$(f,\{\gamma_{e}^{i}\}_{e\in E(\Gamma),i\in[N]})$
where $f:V(\Gamma)\to V(Y)$ is an injective function, and each $\gamma_{e}^{i}$ is a walk in $Y$ such that $\gamma_{e}:=\gamma_{e}^{1}\ast\dots\ast\gamma_{e}^{N}$ (concatenation of walks) is a walk with endpoints given by $f(\text{endpoints}(e))$. Additionally, for any two non-adjacent edges $e,e^{\prime}\in E(\Gamma)$ and any $i\in[N]$, we have:
$\gamma_{e}^{i}\cap\gamma_{e^{\prime}}^{i}=\emptyset.$
In this step we prove the following lemma:
Lemma 3.4.
Let $\Gamma$ be a matching graph, if $f:V(\Gamma)\rightarrow V(P_{0})$ is an injective function, then there is a $(6n-3)$-piecewise embedding $(\tilde{f},\{\gamma_{e}^{i}\}_{e\in E(\Gamma),i\in[6n-3]})$ from $\Gamma$ to $P_{1}$ such that $\tilde{f}v=(0,fv)$ for every $v\in V(\Gamma)$.
Proof.
Define $N:=6n-3$. Let $\Gamma=(V,E)$ be a matching graph, and let $f:V\rightarrow V(P_{0})$ be an injective function. We will find a set $\{\gamma_{e}^{i}\}_{e\in E,i\in[N]}$ such that $(\tilde{f},\{\gamma_{e}^{i}\}_{e\in E,i\in[N]})$ forms a $N$-piecewise embedding of $\Gamma$ into $P_{1}$ where $\tilde{f}v=(0,fv)$ for every $v\in V(\Gamma)$.
To construct the desired set $\{\gamma_{e}^{i}\}_{e\in E,i\in[N]}$, we use proposition 1.2 (box permutation decomposition) by finding a relevant permutation for the problem. Since $\Gamma$ is a matching graph, there is a permutation $\sigma\in\text{Per}(V(P_{0}))=\text{Per}(V_{1}\times...\times V_{n})$ such that if $u\sim v$ then $\sigma(fu)=fv$.
Using proposition 1.2, we obtain a set of permutations $\sigma_{1},...,\sigma_{2n-1}\in\text{Per}(V_{1}\times...\times V_{n})$ such that
$\sigma=\sigma_{1}\circ\dots\circ\sigma_{2n-1}$
where each $\sigma_{i}$ is a one-dimensional permutation. This implies that for every $i\in[2n-1]$, there exists a coordinate $j_{i}$ such that, for every $\vec{v}\in V_{1}\times\dots\times V_{n}$, the vectors $\vec{v}$ and $\sigma_{i}(\vec{v})$ can differ only in the $j_{i}$ coordinate. Since $G_{j_{i}}$ is connected, there exists a path in $G_{1}\Box\dots\Box G_{n}$ from $\vec{v}$ to $\sigma_{i}(\vec{v})$ that remains constant on all coordinates except maybe $j_{i}$. Let us denote this path by $P_{i,\vec{v}}$.
Next, we define $\{\gamma_{e}^{i}\}_{e\in E,i\in[N]}$. First define $\tau_{i}:=\sigma_{1}\circ\dots\circ\sigma_{i}$ for $i\in[2n-1]$ and $\tau_{0}:=id$. For an edge $e\in E$ we choose an endpoint $u\in e$ and define $\{\gamma_{e}^{i}\}_{i\in[k]}$ in trios: For each $i\in[2n-1]$ define the path $\gamma^{3i-2}_{e}$ to be
$(0,\tau_{i-1}(fu))\stackrel{{\scriptstyle 0\rightsquigarrow[\tau_{i-1}(fu)]_{j_{i}}}}{{\rightsquigarrow}}([\tau_{i-1}(fu)]_{j_{i}},\tau_{i-1}(fu)),$
the path $\gamma_{e}^{3i-1}$ to be
$([\tau_{i-1}(fu)]_{j_{i}},\tau_{i-1}(fu))\stackrel{{\scriptstyle P_{i,\tau_{i-1}(fu)}}}{{\rightsquigarrow}}([\tau_{i-1}(fu)]_{j_{i}},\tau_{i}(fu)),$
and the path $\gamma_{e}^{3i}$ to be
$([\tau_{i-1}(fu)]_{j_{i}},\tau_{i}(fu))\stackrel{{\scriptstyle[\tau_{i-1}(fu)]_{j_{i}}\rightsquigarrow 0}}{{\rightsquigarrow}}(0,\tau_{i}(fu)).$
It is straightforward to see that $\{\gamma_{e}^{i}\}_{i\in[N]}$ was chosen so that $\gamma_{e}:=\gamma_{e}^{1}\ast\dots\ast\gamma_{e}^{N}$ forms a walk from $\tilde{f}u$ to $\tilde{f}v$. We now verify that this indeed constitutes an $N$-piecewise embedding:
  1. (1)
    $\tilde{f}$ is injective because $f$ is injective.
  2. (2)
    For any two non-adjacent edges $e,e^{\prime}\in E$, let $u\in e,u^{\prime}\in e^{\prime}$ be the vertices such that $\gamma_{e}^{1}[0]=(0,fu)$ and $\gamma_{e^{\prime}}^{1}[0]=(0,fu^{\prime})$. Let $j\in[2n-1]$; we claim that $\gamma_{e}^{j}\cap\gamma_{e^{\prime}}^{j}=\emptyset$. Indeed, if $j\in 3\mathbb{Z}+\{0,1\}$, it is easy to observe that $\gamma_{e}^{j}\cap\gamma_{e^{\prime}}^{j}=\emptyset$. Now, assume $j\equiv-1\pmod{3}$ and set $i:=(j+1)/3$. If $\gamma_{e}^{j}$ and $\gamma_{e^{\prime}}^{j}$ intersect, then $[\tau_{i-1}(fu)]_{j_{i}}=[\tau_{i-1}(fu^{\prime})]_{j_{i}}$, and the paths $P_{i,\tau_{i-1}(fu)}$ and $P_{i,\tau_{i-1}(fu^{\prime})}$ also intersect. Since every point in $P_{i,\tau_{i-1}(fu)}$ can differ from $\tau_{i-1}(fu)$ only at the coordinate $j_{i}$, and the same is true for $u^{\prime}$, it follows that $\tau_{i-1}(fu)$ and $\tau_{i-1}(fu^{\prime})$ can differ only at $j_{i}$. However, as $\tau_{i-1}(fu)\neq\tau_{i-1}(fu^{\prime})$, we must have $[\tau_{i-1}(fu)]_{j_{i}}\neq[\tau_{i-1}(fu^{\prime})]_{j_{i}}$, leading to a contradiction.
3.5. Step 2: From piecewise embedding to combinatorial embedding
In this step we prove:
Lemma 3.5.
Let $\Gamma$ be a matching graph, if $f:V(\Gamma)\rightarrow\{(0,0)\}\times V(P_{0})$ is an injective function, then there is a combinatorial embedding embedding $\tilde{f}:\Gamma\rightarrow P_{2}$ such that $\tilde{f}|_{V(\Gamma)}=f$.
The following claim together with lemma 3.4 immediately prove this.
Lemma 3.6.
For a graph $\Gamma$ and $N\in\mathbb{N}$, let $(f,\{\gamma_{e}^{i}\}_{e\in E(\Gamma),i\in[N]})$ be a $N$-piecewise embedding of $\Gamma$ into some graph $Y$. There exists a combinatorial embedding $\tilde{f}:G\to C_{N+1}\Box Y$ where $\tilde{f}v=(0,fv)$ for every $v\in V$.
Proof.
Denote $\Gamma=(V,E)$. For $v\in V$, define
$\tilde{f}v:=(0,fv).$
For $e\in E$, define $\tilde{f}_{e}$ as follows:
$(0,\gamma_{e}^{1}[0])$ $\rightarrow(1,\gamma_{e}^{1}[0])\stackrel{{\scriptstyle\gamma_{e}^{1}}}{{\rightsquigarrow}}(1,\gamma_{e}^{1}[-1])$
$\rightarrow(2,\gamma_{e}^{2}[0])\stackrel{{\scriptstyle\gamma_{e}^{2}}}{{\rightsquigarrow}}(2,\gamma_{e}^{2}[-1])$
$\vdots$
$\rightarrow(N,\gamma_{e}^{N}[0])\stackrel{{\scriptstyle\gamma_{e}^{N}}}{{\rightsquigarrow}}(N,\gamma_{e}^{N}[-1])\rightarrow(0,\gamma_{e}^{N}[-1]).$
This is indeed a walk with endpoints $\tilde{f}(\text{endpoints}(e))$. Now we verify that $\tilde{f}$ is a combinatorial embedding:
  1. (1)
    $\tilde{f}|_{V}$ is injective since $f|_{V}$ is injective.
  2. (2)
    Let $e,e^{\prime}\in E$ be two non-adjacent edges. Assume there is $(\xi_{1},\xi_{2})\in\tilde{f}_{e}\cap\tilde{f}_{e^{\prime}}$, so $\xi_{1}=i$ for some $0\leq i\leq k$.
    • If $i=0$, then $\xi_{2}\in f(\text{endpoints}(e))\cap f(\text{endpoints}(e^{\prime}))=\emptyset$, a contradiction.
    • If $1\leq i\leq N$, then $\xi_{2}\in\gamma^{i}_{e}\cap\gamma^{i}_{e^{\prime}}=\emptyset$, another contradiction.
  3. (3)
    Let $e$ be an edge and $v$ a vertex not in it. Since $\tilde{f}v=(0,fv)$, if $\tilde{f}v\in\tilde{f}_{e}$, then $fv\in f(\text{endpoints}(e))$, which is impossible as $f|_{V}$ is injective.
3.6. Step 3: From matching graph to graph with maximal degree $\leq 3$
In this step we conclude theorem 3.2:
Proof.
Let $\Gamma$ be a graph with maximal degree $\leq 3$ and $\leq\frac{1}{4}|V_{1}|\cdots|V_{n}|$ vertices. Denote by $\Gamma^{\prime}=(V,E)$ its simple subdivision, we will construct a combinatorial embedding of $\Gamma^{\prime}$ into $P_{3}$ which will imply $\Gamma\leq P_{3}$, as needed.
Notice that $|V|\leq 4|V(\Gamma)|\leq|V_{1}|\cdots|V_{n}|$, so there is an injective function $f:V\rightarrow\{(0,0)\}\times V_{1}\times\ldots\times V_{n}$. By Vizing’s Theorem, we can partition $E$ into $4$ disjoint subsets $E_{0},\dots,E_{3}$ such that for each $i$, the graph $(V,E_{i})$ is a matching graph. By lemma 3.5, for any $i\in\{0,1,2,3\}$, there is a combinatorial embedding $f^{(i)}$ of $(V,E_{i})$ into $P_{2}$ that extends $f$.
We construct a combinatorial embedding $\tilde{f}:\Gamma^{\prime}\rightarrow P_{3}$: For $v\in V$, define $\tilde{f}v:=(0,fv)$. For $i\in\{0,1,2,3\}$ and $e=uv\in E_{i}$, define $\tilde{f}_{e}$ as
$(0,fu)\stackrel{{\scriptstyle 0\rightsquigarrow i}}{{\rightsquigarrow}}(i,fu)\stackrel{{\scriptstyle f^{(i)}_{e}}}{{\rightsquigarrow}}(i,fv)\stackrel{{\scriptstyle i\rightsquigarrow 0}}{{\rightsquigarrow}}(0,fv).$
Now we verify that $\tilde{f}$ is a combinatorial embedding:
  • $\tilde{f}|_{V}$ is injective since $f|_{V}$ is injective.
  • Let $e,e^{\prime}\in E$ be two non-adjacent edges, and let $i$ and $i^{\prime}$ be the indices such that $e\in E_{i}$ and $e^{\prime}\in E_{i^{\prime}}$. We aim to show that $\tilde{f}_{e}\cap\tilde{f}_{e^{\prime}}=\emptyset$. By projecting these walks onto the second component, we obtain $f^{(i)}_{e}$ and $f^{(i^{\prime})}_{e^{\prime}}$. Since $f_{e}^{(i)}\cap f(\text{endpoints}(e^{\prime}))=\emptyset$ and vice versa, we conclude that any intersection between $\tilde{f}_{e}$ and $\tilde{f}_{e^{\prime}}$ could only happen in the ”middle phase” (the second arrow in the definition of $\tilde{f}_{e}$). If $i\neq i^{\prime}$, no intersection can occur. If $i=i^{\prime}$, an intersection is also impossible because $f^{(i)}$ is a combinatorial embedding, so $f^{(i)}_{e}\cap f^{(i)}_{e^{\prime}}=\emptyset$.
  • Let $e$ be an edge and $v$ a vertex not in $e$. There is $i$ such that $e\in E_{i}$. We need to show that $\tilde{f}v\notin\tilde{f}_{e}$, which follows immediately from $fv\notin f^{(i)}_{e}$.
4. Box permutation decomposition
Here we prove proposition 1.2 about decomposition of permutation of a box. We will first prove a claim called the well-dispersed lemma which will help us prove the upper bound of the proposition (that any permutation can be decomposed into $12d-$ 1 one-dimensional permutations). We finish this section by proving the lower bound of the proposition (that there exists a permutation such that any decomposition into one-dimensional permutations requires $12d-$ 1 elements).
Definition 4.1.
Let $d\in\mathbb{N}$ and $n_{1},...,n_{d}\in\mathbb{N}$. Define $X:=[n_{1}]\times...\times[n_{d}]$. A permutation $\sigma:X\rightarrow X$ is called $\bm{i}$-permutation if it affects only the $i$-coordinate.
4.1. Well-dispersed Lemma
Definition 4.2.
A matrix $A$ is well-dispersed if each column is a vector with all distinct components.
Lemma 4.3 (well-dispersed Lemma).
Let $L$ be a finite set of labels, and let $m:=|L|$ be its size. Suppose $A\in M_{m\times n}(L)$ is a matrix such that each label $\ell\in L$ appears in exactly $n$ entries of $A$. Then, there exist $m$ permutations $\sigma_{1},\dots,\sigma_{m}\in S_{n}$ such that the matrix $\tilde{A}:=(A_{i\sigma_{i}(j)})_{i,j}$ is well-dispersed.
$A\hfil\hskip 56.9055pt&\tilde{A}\\ \begin{array}[]{|c|c|c|c|}\hline\cr 1&3&5&5\\ \hline\cr 6&4&2&2\\ \hline\cr 5&2&1&3\\ \hline\cr 4&1&6&3\\ \hline\cr 1&2&4&3\\ \hline\cr 4&6&6&5\\ \hline\cr\end{array}\hfil\hskip 56.9055pt&\begin{array}[]{|c|c|c|c|}\hline\cr 1&5&5&3\\ \hline\cr 6&4&2&2\\ \hline\cr 5&2&3&1\\ \hline\cr 3&1&4&6\\ \hline\cr 2&3&1&4\\ \hline\cr 4&6&6&5\\ \hline\cr\end{array}\end{array}$
Figure 2. Well-dispersed lemma example.
Proof.
Observe that since the label set $L$ has size $m$, a matrix in $M_{m\times n}(L)$ is well-dispersed if and only if each column contains all the labels.
We prove the statement by induction on the number of columns. When there is only one column, each label appears exactly once, making $A$ well-dispersed. Now, assume $n>1$:
If the first column of $A$ contains all the labels, the result follows by the induction hypothesis. Otherwise, assume there exists a proper subset $L^{\prime}\subsetneq L$, representing the set of all labels in the first column of $A$. It suffices to find permutations $\sigma_{1},\dots,\sigma_{m}\in S_{n}$ such that the first column of the matrix $(A_{i\sigma_{i}(j)})_{i,j}$ will contain $|L^{\prime}|+1$ labels.
Since the first column of $A$ does not contain all the labels, there must be a label $\ell_{0}\in L$ that appears twice in this column. Let $I_{0}:=\{i_{0},i_{1}\}$ denote two indices such that $A_{i_{0}1}=A_{i_{1}1}=\ell_{0}$, and define $L_{0}:=\{A_{ij}\mid(i,j)\in I_{0}\times[n]\}$ to be the set of labels in the rows indexed by $I_{0}$. For $k\in[m]$, define
$I_{k}:=\{i\in[m]\mid A_{i1}\in L_{k-1}\}$
$L_{k}:=\{A_{ij}\mid(i,j)\in I_{k}\times[n]\}$
Notice that $I_{k}\subseteq I_{k+1}$ and $L_{k}\subseteq L_{k+1}$.
We claim that each label in $L_{m}$ can be reached. Specifically, for every $(a,b)\in I_{m}\times[n]$, there exists a way to permute elements within each row of $A$ such that the resulting matrix has labels $L^{\prime}\cup\{A_{ab}\}$ in the first column. Given this, it remains to show that $L_{m}$ contains at least one label from $L\setminus L^{\prime}$. These claims will be proven in the lemmas below. ∎
Lemma 4.4.
For every $0\leq k\leq m$ and every entry $(a,b)\in I_{k}\times[n]$, there exist permutations $\{\tau_{i}\in S_{n}\}_{i\in I_{k}}$ such that $\{A_{i\tau_{i}(1)}\}_{i\in I_{k}}=\{A_{i1}\}_{i\in I_{k}}\cup\{A_{ab}\}$. In other words, after applying these permutations to the rows indexed by $I_{k}$, the new first column will contain the label $A_{ab}$ without losing any existing labels.
Proof.
We prove this by induction on $k$. For the base case, let $k=0$ and $(a,b)\in\{i_{0},i_{1}\}\times[n]$. Without loss of generality, assume $a=i_{0}$. Then, set $\tau_{i_{0}}:=(1b)$ and $\tau_{i_{1}}:=\text{id}$. Indeed, we have $\{A_{i_{0}\tau_{i_{0}}(1)},A_{i_{1}\tau_{i_{1}}(1)}\}=\{A_{i_{0}b},A_{i_{1}1}\}=\{A_{i_{0}1},A_{i_{1}1}\}\cup\{A_{ab}\}$.
Now, consider the case $k>0$ and $(a,b)\in I_{k}\times[n]$. If $a\in I_{k-1}$, we conclude by the induction hypothesis. Otherwise, assume $a\in I_{k}\setminus I_{k-1}$. Our goal is to swap the $1$st and $b$-th components of the $a$-th row so that the label $A_{ab}$ appears in the first column. To achieve this without losing the label $A_{a1}$, we use the induction hypothesis to permute the components in the rows indexed by $I_{k-1}$so that the first column contains $A_{a1}$ twice!
[…]
Figure 3. Well-dispersed lemma proof strategy.
By the definition of $I_{k}$, there exist $a^{\prime}\in I_{k-1}$ and $b^{\prime}\in[n]$ such that $A_{a1}=A_{a^{\prime}b^{\prime}}$. By the induction hypothesis, there are permutations $\{\tau_{i}\in S_{n}\}_{i\in I_{k-1}}$ such that
$\{A_{i\tau_{i}(1)}\}_{i\in I_{k-1}}=\{A_{i1}\}_{i\in I_{k-1}}\cup\{A_{a^{\prime}b^{\prime}}\}.$
Now, define $\tau_{a}:=(1b)$, and for $i\in I_{k}\setminus(I_{k-1}\cup\{a\})$, set $\tau_{i}:=\text{id}$. This construction gives the desired result, as follows:
$\{A_{i\tau_{i}(1)}\}_{i\in I_{k}}$ =$=\{A_{i\tau_{i}(1)}\}_{i\in I_{k-1}}\cup\{A_{i\tau_{i}(1)}\}_{i\in I_{k}\setminus(I_{k-1}\cup\{a\})}\cup\{A_{a\tau_{a}(1)}\}$
=${=}\{A_{i1}\}_{i\in I_{k-1}}\cup\{A_{a^{\prime}b^{\prime}}\}\cup\{A_{i1}\}_{i\in I_{k}\setminus(I_{k-1}\cup\{a\})}\cup\{A_{ab}\}$
=${=}\{A_{i1}\}_{i\in I_{k}}\cup\{A_{ab}\}.$
Lemma 4.5.
$L_{m}\not\subseteq L^{\prime}$.
Proof.
We will prove by induction on $k\in[m]$ that if $L_{k-1}\subseteq L^{\prime}$, then $|I_{k}|>k$.
For the base case $k=1$, we have $|I_{1}|\geq|I_{0}|>1$. For $k>1$, assume that $L_{k-1}\subseteq L^{\prime}$. Then $L_{k-2}\subseteq L^{\prime}$ as well. By the induction hypothesis, we have $|I_{k-1}|\geq k$. The set $L_{k-1}$ consists of labels of $|I_{k-1}|\cdot n$ entries of $A$, with each label appearing at most $n$ times, so $|L_{k-1}|\geq|I_{k-1}|\geq k$. Since $L_{k-1}\subseteq L^{\prime}$, each label in $L_{k-1}$ appears in $\{A_{i1}\}_{i\in I_{k}}$. Therefore, |$|\{A_{i1}\}_{i\in I_{k}}|\geq k$. However, since $A_{i_{0}1}=A_{i_{1}1}$, we must have $|I_{k}|>k$. ∎
4.2. Upper bound - box permutation decomposition
Now are ready to prove the upper bound of proposition 1.2:
Proposition 4.6 (upper bound of proposition 1.2).
Let $n_{1},\dots,n_{d}\in\mathbb{N}$ and define $X:=[n_{1}]\times\dots\times[n_{d}]$. Every permutation $\sigma:X\to X$ can be written as
$\sigma=\sigma_{d}\circ\dots\circ\sigma_{2}\circ\sigma_{1}\circ\sigma^{\prime}_{2}\circ\dots\circ\sigma^{\prime}_{d},$
where $\sigma_{i}$ and $\sigma^{\prime}_{i}$ are $i$-permutations.
Proof.
The case $d=1$ is trivial, in this case $\sigma$ is already a $1$-permutation. We will prove only the case $d=2$ as it implies the case $d>2$: $\sigma$ is a permutation of $[n_{1}]\times...\times[n_{d}]$, it can be also seen as permutation of $Y\times[n_{d}]$ for $Y=[n_{1}]\times...\times[n_{d-1}]$. Hence, given the case of $d=2$ we know that $\sigma=\sigma_{d}\circ\tau\circ\sigma_{d}^{\prime}$ where $\sigma_{d},\sigma_{d}^{\prime}$ are $d$-permutations and $\tau$ is a permutation that preserve the $d$-coordinate. Hence for every $i\in[n_{d}]$, the function $\tau_{i}:=\tau|_{Y\times\{i\}}$ is a permutation. So, by induction, there is a decomposition
$\tau_{i}=\sigma_{d-1,i}\circ...\circ\sigma_{2,i}\circ\sigma_{1,i}\circ\sigma^{\prime}_{2,i}\circ...\circ\sigma^{\prime}_{d-1,i}$
where $\sigma_{j,i},\sigma^{\prime}_{j,i}$ are $j$-permutations in $Y\times\{i\}\rightarrow Y\times\{i\}$ for every $j\in[d-1]$.
Fix $j\in[d-1]$, we have the $j$-permutations $\sigma_{j,1},...,\sigma_{j,n_{d}}$ of $Y\times\{1\},...,Y\times\{n_{d}\}$ respectively, we want to unify all of them to be a one $j$-permutation of $X$, hence we define $\sigma_{j}:X\rightarrow X$ as $x\mapsto\sigma_{j,x_{d}}(x)$, and similarly define $\sigma_{j}^{\prime}$. Since $\tau(x)=\tau_{x_{d}}(x)$ we get $\tau=\sigma_{d-1}\circ...\circ\sigma_{2}\circ\sigma_{1}\circ\sigma^{\prime}_{2}\circ...\circ\sigma^{\prime}_{d-1}$ as needed.
Now we prove the $d=2$ case, so assume $X=[m]\times[n]$. Define the matrix $A:=(\sigma^{-1}(i,j))_{ij}$, notice that we need to find $2$-permutations $\sigma_{2},\sigma_{2}^{\prime}\in Per(X)$ and $1$-permutation $\sigma_{1}\in Per(X)$ such that $A_{\sigma_{2}\circ\sigma_{1}\circ\sigma_{2}^{\prime}(i,j)}=(i,j)$ for every $(i,j)\in[m]\times[n]$. Since each number $i\in[m]$ appears as the first component in exactly $n$ entries of $A$, we can use the Well-dispersed lemma to permute the elements in each row so that every column contains all possible first components. After this, we can rearrange the elements within each column so that the first components of the columns are in the order $(1,\dots,m)$. At this stage, for each $i\in[m]$, the labels in the $i$-th row all start with the first component $i$. Finally, we can permute the elements within each row so that the $i$-th row becomes $(i,1),\dots,(i,n)$, completing the proof.
$\begin{array}[]{|c|c|c|}\hline\cr 12&42&13\\ \hline\cr 11&22&33\\ \hline\cr 21&31&41\\ \hline\cr 32&23&43\\ \hline\cr\end{array}$ $A$$\begin{array}[]{|c|c|c|}\hline\cr 42&12&13\\ \hline\cr 11&22&33\\ \hline\cr 21&31&41\\ \hline\cr 32&43&23\\ \hline\cr\end{array}$
Well-dispersed
lemma
Sorting first
components in
columns
Sorting second
components in
rows
[…]
Figure 4. $d=2$ strategy.
4.3. Lower bound: box permutation decomposition
Proposition 4.7 (lower bound of proposition 1.2).
Let $d\in\mathbb{N}$ and $n_{1},\ldots,n_{d}\geq 2$. Define $X:=[n_{1}]\times\cdots\times[n_{d}]$. There exists a permutation $\sigma:X\to X$ such that, for any expression of $\sigma$ as a composition $\sigma=\sigma_{1}\circ\cdots\circ\sigma_{r}$ where each $\sigma_{i}$ is a one-dimensional permutation, we have $r\geq 2d-1$.
Proof.
Let $\sigma$ be the permutation that swaps $(1,\ldots,1)$ and $(2,\ldots,2)$. Suppose $\sigma=\sigma_{1}\circ\cdots\circ\sigma_{r}$, where each $\sigma_{i}$ is a non-identity one-dimensional permutation. Since $\sigma$ swaps $(1,\ldots,1)$ and $(2,\ldots,2)$, there must be at least one $i$-permutation in this decomposition for every $i\in[d]$. If for each $i\in[d]$ there are at least two $i$-permutations in the decomposition, then there are at least $2d$ permutations, and we are done. Assume instead that there exists $i\in[d]$ such that there is exactly one $i$-permutation in the decomposition, say $\sigma_{j}$.
Define $\tau^{\prime}=\sigma_{1}\circ\cdots\circ\sigma_{j-1}$ and $\tau=\sigma_{j+1}\circ\cdots\circ\sigma_{r}$. The permutation $\sigma_{j}$ is the $i$-permutation that swaps $\tau(1,\ldots,1)$ and $\tau(2,\ldots,2)$. Since, for any $p\in X\setminus\{(1,\ldots,1),(2,\ldots,2)\}$, we have $\sigma_{j}(\tau(p))=\tau(p)$. This follows because $[\sigma_{j}(\tau(p))]_{i}=[\tau^{\prime}(\sigma_{j}(\tau(p)))]_{i}=[\sigma(p)]_{i}=p_{i}=[\tau(p)]_{i}.$
Since $\sigma_{j}$ swaps $\tau(1,\ldots,1)$ and $\tau(2,\ldots,2)$ and is an $i$-permutation, $\tau(1,\ldots,1)$ and $\tau(2,\ldots,2)$ must agree in $d-1$ coordinates. However, $(1,\ldots,1)$ and $(2,\ldots,2)$ disagree in all coordinates. Thus, $\tau$ must consist of at least $d-1$ one-dimensional permutations. By the same reasoning, $\tau^{\prime}$ also consists of at least $d-1$ one-dimensional permutations. Consequently, the original decomposition consists of at least $12(d-1)+1=2d-$ 1 one-dimensional permutations.
5. Upper bound: minor-universality
We begin by proving the upper bound of theorem A, as its proof captures the core idea of the generalized version while being simpler and more elegant. In the next subsection, we will address the generalized version, corresponding to the second part of theorem B.
Theorem 5.1 (upper bound of theorem A).
There is an absolute constant $C$ such that $Q_{d}$ is not $\frac{C2^{d}}{\sqrt{d}}$-minor-universal.
The idea is to utilize an expander graph as one that is hard to embed. This difficulty arises from examining how easily $Q_{d}$ can be separated, drawing inspiration from the concept of the separation profile introduced in [BST12].
Since we will use expander graphs to prove this theorem, we provide here the precise definition of the Cheeger constant as follows:
Definition 5.2.
Let $G=(V,E)$ be a graph, and let $U,W\subseteq V$ be disjoint subsets. We define $\mathcal{E}(U,W):=\{e\in E\mid|e\cap U|=1,|e\cap W|=1\}$. The Cheeger constant of $G$ is
$h(G):=\inf_{\emptyset\neq U\subseteq V,|U|\leq\frac{1}{2}|V|}\frac{|\mathcal{E}(U,U^{C})|}{|U|}.$
Proof.
Let $\{G_{i}\}_{i\in 2\mathbb{N}}$ be a family of expander graphs (without isolated vertices) with Cheeger constant at least $h$ (for some constant $0<h\leq 1$) and maximal degree at most $3$, such that $|V(G_{i})|=i$.  [^4] It is well known that such a family exists, see e.g. [Ko24, Theorem 4.1.1].
Define the constant $C:=\frac{10\sqrt{2}e}{\sqrt{\pi}h}$, let $d\in\mathbb{N}$ and choose some even number $n$ such that $C\frac{2^{d}}{\sqrt{d}}\leq n\leq 2C\frac{2^{d}}{\sqrt{d}}$. Consider the graph $G:=G_{n}$ and write $G=(V,E)$. This graph has at most $1.5n\leq 3C\frac{2^{d}}{\sqrt{d}}$ edges and it is not a minor of $Q_{d}$, as needed. We will show that $G\not\leq Q_{d}$ by showing that if $G\leq Q_{d}$, then $C\frac{2^{d}}{\sqrt{d}}>n$, which is a contradiction.
Assume that $G\leq Q_{d}$. Then, by lemma 2.2, we have a combinatorial embedding $f:G\rightarrow Q_{d}$. We will partition $V(Q_{d})=\{0,1\}^{d}$ into two parts using a relatively small cut $S$ such that each part contains approximately half of the vertices of $f(V)$. Since $G$ has a large Cheeger constant, there are many roads of $f$ that connect vertices of $f(V)$ from one part, to vertices of $f(V)$ from the other part. A large subset of these roads is pairwise disjoint, meaning that each road must pass through a distinct point in $S$, which implies that $S$ is large. Since $S$ is small relative to the cube $Q_{d}$, it follows that the cube itself must be large, which means that $\frac{2^{d}}{\sqrt{d}}$ is large.
We will find this $S$ by considering all the spheres around the point $\vec{0}$. For $0\leq k\leq d$, let $B_{k}$ be the open ball in $Q_{d}$ of radius $k$ around $\vec{0}$, and similarly let $S_{k}$ denote the corresponding sphere of radius $k$. Note that
$|S_{k}|={d\choose k}\leq{d\choose\lceil\frac{d}{2}\rceil}<\frac{\sqrt{2}e}{\sqrt{\pi}}\frac{2^{d}}{\sqrt{d}}.$
If there exists $k$ such that $|S_{k}|>0.1n$, then $n<\frac{10\sqrt{2}e}{\sqrt{\pi}}\frac{2^{d}}{\sqrt{d}}\leq C\frac{2^{d}}{\sqrt{d}}$ and we finish. Therefore, we assume $|S_{k}|\leq 0.1n$ for all $k$.
For a subset $A\subseteq\mathbb{Z}_{2}^{d}$, define $\rho(A):=|(f|_{V})^{-1}(A)|$, which represents the number of vertices that $f$ maps to $A$. We obtain the following values:
$\rho(B_{0})=\rho(\emptyset)=0,$
$\rho(B_{d})\geq 0.9n,$
$\rho(B_{k+1})\leq\rho(B_{k})+0.1n.$
Thus, there exists $0<r<d$ such that $0.4n\leq\rho(B_{r})\leq 0.5n$. Define $U:=(f|_{V})^{-1}(B_{r})$, then $0.4n\leq|U|\leq 0.5n$, and hence $|\mathcal{E}(U,U^{C})|\geq h|U|\geq 0.4hn$. For each edge $e\in\mathcal{E}(U,U^{C})$, the road $f_{e}$ passes through $S_{r}$. Since $G$ is a graph with maximal degree at most 3, Vizing’s Theorem tells us that there exists a subset of $\mathcal{E}(U,U^{C})$ of size at least $\frac{1}{4}|\mathcal{E}(U,U^{C})|$ in which every pair of edges is non-adjacent. Thus all the associated roads are pairwise disjoint and pass through $S_{r}$. Therefore, $|S_{r}|\geq\frac{1}{4}|\mathcal{E}(U,U^{C})|\geq 0.1hn$.
Thus, we have
$n\leq\frac{10}{h}|S_{r}|<\frac{10\sqrt{2}e}{\sqrt{\pi}h}\frac{2^{d}}{\sqrt{d}}\leq C\frac{2^{d}}{\sqrt{d}}.$
5.1. Generalized upper bound
Theorem 5.3 (upper bound of theorem B).
Let $2\leq k\in\mathbb{N}$, there exists a constant $C=C(k)$ that satisfies the following: Let $n\in\mathbb{N}$ and let $G_{1},\ldots,G_{n}$ be sequence of graphs, each with $\leq k$ vertices. The graph $G_{1}\Box\ldots\Box G_{n}$ is not $\frac{Ck^{n}}{\sqrt{n}}$-minor-universal.
Proof.
We rely on the proof of the previous theorem; the difference in the current proof lies in the need for slightly more computations.
To simplify the computations, we replace $k$ with $k+1$. Thus, we now assume $k\geq 1$, each graph $G_{i}$ has $\leq k+1$ vertices, and we aim to construct a graph with $\leq\frac{C(k+1)^{n}}{\sqrt{n}}$ edges that is not a minor of $G_{1}\Box\cdots\Box G_{n}$.
Consider the graph $G:=\underbrace{K_{k+1}\Box\cdots\Box K_{k+1}}_{\text{n times}}$ where $K_{k+1}$ is the complete graph. Since $G_{1}\Box\cdots\Box G_{n}\leq G$, finding a graph with $\leq\frac{C(k+1)^{n}}{\sqrt{n}}$ edges and with no isolated vertices, that is not a minor of $G$, ensures that it is also not a minor of $G_{1}\Box\cdots\Box G_{n}$, which is our goal. This is exactly what we are going to do.
We label the vertices of each $K_{k+1}$ by $\{0,\ldots,k\}$, so the vertices of $G$ correspond to strings in $\{0,\ldots,k\}^{n}$. Referring to the proof of theorem 5.1, it suffices to bound the sizes of spheres in $G$. Specifically, we need to show the existence of a constant $C=C(k)$ such that every sphere in $G$ centered at $\vec{0}$ has size $\leq\frac{C(k+1)^{n}}{\sqrt{n}}$.
Let $0\leq r\leq n$. A sphere of radius $r$ around $\vec{0}$ has size $\binom{n}{r}k^{r}$. We need to show that there exists a constant $C=C(k)$ such that $\binom{n}{r}k^{r}\leq\frac{C(k+1)^{n}}{\sqrt{n}}$ for all $0\leq r\leq n$.
The easy cases $r\in\{0,n\}$ will be handled separately. For now, fix $11\leq r\leq n-$ 1 and let $a=\frac{r}{n}$. Using the inequality
$\sqrt{2\pi n}\left(\frac{n}{e}\right)^{n}e^{\frac{1}{12n+1}}<n!<\sqrt{2\pi n}\left(\frac{n}{e}\right)^{n}e^{\frac{1}{12n}},$
(from [R55]), we obtain
$\binom{n}{an}k^{an}\leq\frac{c}{\sqrt{a(1-a)n}}\left(\frac{k^{a}}{a^{a}(1-a)^{1-a}}\right)^{n},$
where $c=\frac{e}{\sqrt{2\pi}}$. Define $f(a):=\frac{k^{a}}{a^{a}(1-a)^{1-a}}$. We then need to prove the existence of a constant $C=C(k)$ such that
$\frac{cf(a)^{n}}{\sqrt{a(1-a)n}}\leq\frac{C(k+1)^{n}}{\sqrt{n}}$
for every $a\in\left\{\frac{1}{n},\ldots,\frac{n-1}{n}\right\}$.
By differentiating $f$ and comparing it to 00, one can conclude that: for $x\in(0,1)$, we have $f(x)\leq k+1$, with equality $f(x)=k+1$ if and only if $x=\frac{k}{k+1}$. Moreover, $f^{\prime}(x)>0$ on $(0,\frac{k}{k+1})$ and $f^{\prime}(x)<0$ on $(\frac{k}{k+1},1)$.
We proceed by selecting a value $\frac{k}{k+1}<k^{\prime}<1$ and defining $m:=\sup_{a\in(0,1-k^{\prime}]\cup[k^{\prime},1)}f(a)$. Note that $m=\max(f(k^{\prime}),f(1-k^{\prime}))<k+1$, and notice it is just a constant, independent of $n$.
We now derive the desired inequalities:
  1. (1)
    For $a\in\{0,1\}$, we have
    $\binom{n}{an}k^{an}\leq k^{n}\leq C_{1}\frac{(k+1)^{n}}{\sqrt{n}},$
    for some constant $C_{1}=C_{1}(k)$.
  2. (2)
    For $a\in\{\frac{1}{n},\ldots,\frac{n-1}{n}\}\cap((0,1-k^{\prime}]\cup[k^{\prime},1))$, we have
    $\binom{n}{an}k^{an}\leq\frac{cf(a)^{n}}{\sqrt{a(1-a)n}}\leq\frac{cm^{n}}{\sqrt{\frac{n(n-1)}{n^{2}}}}\leq C_{2}\frac{(k+1)^{n}}{\sqrt{n}},$
    for some constant $C_{2}=C_{2}(k)$.
  3. (3)
    For $a\in\{\frac{1}{n},\ldots,\frac{n-1}{n}\}\cap[1-k^{\prime},k^{\prime}]$, we have
    $\binom{n}{an}k^{an}\leq\frac{cf(a)^{n}}{\sqrt{a(1-a)n}}\leq\frac{c}{k^{\prime}(1-k^{\prime})}\frac{(k+1)^{n}}{\sqrt{n}}.$
This completes the proof as required. ∎